Simplify (a’ * b’ * c) + (a * b’ * c) + (a * b’ * c’)
1
Expert's answer
2020-09-28T09:03:24-0400

"\left(\neg \:A\wedge \neg \:B\wedge \:C\right)+\left(A\wedge \neg \:B\wedge \:C\right)+\left(A\wedge \neg \:B\wedge \neg \:C\right)=\left(\neg \:B\wedge \:C\wedge \:A\right)\vee \left(\neg \:B\wedge \:C\wedge \neg \:A\right)\vee \left(A\wedge \neg \:B\wedge \neg \:C\right)=\left(\neg \:B\wedge \:C\wedge \:A\right)\vee \left(\neg \:B\wedge \:C\wedge \neg \:A\right)\vee \left(A\wedge \neg \:B\wedge \neg \:C\right)=\left(\left(\neg \:B\wedge \:C\right)\vee \:A\right)\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:B\right)\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:C\right)=\left(\left(\neg \:B\wedge \:C\right)\vee \:A\right)\wedge \left(\neg \:B\vee \left(\neg \:B\wedge \:C\right)\right)\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:C\right)=\left(\left(\neg \:B\wedge \:C\right)\vee \:A\right)\wedge \left(\neg \:B\vee \left(\neg \:B\wedge \:C\right)\right)\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:C\right)=\left(A\vee \left(\neg \:B\wedge \:C\right)\right)\wedge \neg \:B\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:C\right)=\left(A\vee \left(\neg \:B\wedge \:C\right)\right)\wedge \neg \:B\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:C\right)=\left(A\vee \neg \:B\right)\wedge \left(A\vee \:C\right)\wedge \neg \:B\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:C\right)=\left(A\vee \neg \:B\right)\wedge \left(A\vee \:C\right)\wedge \neg \:B\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:C\right)=\left(A\vee \neg \:B\right)\wedge \left(A\vee \:C\right)\wedge \neg \:B\wedge \left(\left(\neg \:B\wedge \:C\right)\vee \neg \:C\right)=\neg \:B\wedge \left(A\vee \:C\right)\wedge \left(\neg \:C\vee \left(\neg \:B\wedge \:C\right)\right)=\neg \:B\wedge \left(A\vee \:C\right)\wedge \left(\left(\neg \:C\vee \neg \:B\right)\wedge \left(\neg \:C\vee \:C\right)\right)=\neg \:B\wedge \left(A\vee \:C\right)\wedge \left(\left(\neg \:C\vee \neg \:B\right)\wedge \left(\neg \:C\vee \:C\right)\right)=\neg \:B\wedge \left(A\vee \:C\right)\wedge \left(\neg \:C\vee \neg \:B\right)=\neg \:B\wedge \left(\neg \:B\vee \neg \:C\right)\wedge \left(A\vee \:C\right)=\neg \:B\wedge \left(A\vee \:C\right)"